OpenStudy (anonymous):

The area of a rectangle is 72ft^2, and the length of the rectangle is 10 ft less than three times the width. find the dimensions of the rectangle.

OpenStudy (dangerousjesse):

\(\huge\color{darkblue}{w\times(3w-10)=72^2}\)

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

better yet what do I even do lol

OpenStudy (dangerousjesse):

One sec, I need to work through it :P

OpenStudy (anonymous):

okay!

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

have you done quadratics factoring yet?

OpenStudy (triciaal):

the area is 72 do not square it.

OpenStudy (anonymous):

Nope and if I have I don't remember lol

OpenStudy (jdoe0001):

h hmmm ohh yea.. is just 72.. .sorta misread there

OpenStudy (triciaal):

ft^2 is the unit of measure

OpenStudy (jdoe0001):

@atlantaslast well...this one is a quadratic....so... you may want to cover that first

OpenStudy (jdoe0001):